The brakes of a car moving at 14 m/s are applied, and the car comes to a stop in 48 seconds. How far does the car go while it is slowing down from 14 to 10 m/s?


A.13.7 m


B.16.2 m


C.19.5 m


D.26.3 m

1
Expert's answer
2022-05-13T03:24:36-0400

The acceleration of the car

"a=(0-v_i)/t=-14/4.8=-2.92\:\rm m/s^2"

Distance traveled by the car


"d=\frac{v_f^2-v_i^2}{2a}=\frac{10^2-14^2}{2*(-2.92)}=16\; \rm m"

Answer: B